# Amortized Inference of Causal Models via Conditional Fixed-Point Iterations

### v3

@@ Appendix I Experiments on Real World Benchmark
Appendix K Limitations of Cond-FiP
K.1 Evaluating Generalization of Cond-Fip to Larger Sample Size
@@
K.2 Counterfactual Generation with Cond-FiP
We provide results (Table 30) for bechmarking Cond-FiP against baselines for the task of counterfactual generation. We operate in the same setup as the one in our main results ($n_{\mathcal{D_{\mathrm{test}}}}=400$Appendix C and all the methods are provided with the true casual graph. We observe that Unlike the tasks of noise prediction, sample & interventional generation, we find that Cond-FiP is worse than the baselines for the task of counterfactual generation. This can be explained as the training of Cond-FiP decoder relies on the true noise variables, and the model struggles to generalize the learned functional mechanisms when provided with inferred noise variables. We leave the improvement of Cond-FiP for counterfactual generation as future work.

### v4

@@ Appendix H Comparing Cond-FiP with CausalNF
Appendix I Limitations of Cond-FiP
I.1 Evaluating Generalization of Cond-Fip to Larger Sample Size
@@
I.2 Counterfactual Generation with Cond-FiP
We provide results (Table 33) for bechmarking Cond-FiP against baselines for the task of counterfactual generation. We operate in the same setup as the one in our main results ($n_{\mathcal{D_{\mathrm{test}}}}=400$Appendix˜C and all the methods are provided with the true casual graph. We observe that Unlike the tasks of noise prediction, sample & interventional generation, we find that Cond-FiP is worse than the baselines for the task of counterfactual generation. This can be explained as the training of Cond-FiP decoder relies on the true noise variables, and the model struggles to generalize the learned functional mechanisms when provided with inferred noise variables. We leave the improvement of Cond-FiP for counterfactual generation as future work.